OpenStudy (anonymous):

y = ln( x^3 - 5 x + 13) (dy)/(dx) =

OpenStudy (anonymous):

use the chain rule

OpenStudy (anonymous):

any help?

OpenStudy (anonymous):

d/dx[ln y] = y'/y

OpenStudy (anonymous):

another way of doing this is to let u = x^3 - 5x + 13 so y = ln u and dy/du = 1/u and du/dx = 3x^2 - 5 so dy/dx = dy/du * du/dx = 1/u * 3x^2 - 5 = ( 3x^2 - 5) / (x^3 -5x + 13)

OpenStudy (anonymous):

thanx